# Shortcomings of LLMs for Low-Resource Translation: Retrieval and Understanding are Both the Problem

### v1

@@ 6 Discussion
We observe large differences between LLMs, both in terms of the overall quality of their generated translations as well as the effects of prompt type on their outputs. GPT-4o and Gemini, which have the highest baseline scores, benefit least from additional information— their performance with corpus and grammar information actually decreases. This occurs even with manually curated prompts, suggesting it is not an effect of including irrelevant material. On the other hand, it does not represent a ceiling on quality, since both models continue to make errors (GPT-4o produces 10 low-quality translations in our set of 50). These results suggest that even relevant grammar explanations, when written in prose with examples, do little to help the newest generation of LLMs to translate a low-resource language such as Southern Quechua.
@@
Finally, we identify a number of translation errors of varying types that appear to be due to language-specific characteristics, for example ambiguity from syncretism in grammatical markers, polysemous lexical items, or the orthographic and lexical variation discussed in Section 4.2.5. It may be possible to moderate such effects with additional refinement of the retrieval database and methods, which we leave for future work.
@@ 6.1 Ethical concerns
Second, some mistranslations identified in our study appear to draw on stereotypes of indigenous groups (Appendix 6). These are most apparent for the baseline system and GPT-3.5, but also (less frequently) occur with more informative prompts and better LLMs. Stereotypical sentences can involve flowery language with an emphasis on tradition or connectedness to nature Erhart and Hall (2019), as well as the unprompted addition of indigenous Andean cultural customs and products (traditional medicine, chicha) to translations that are otherwise faithful to the source text. The overall effect is to exoticize Southern Quechua speakers and writers in ways that the original sentences do not. Similar stereotypes have also been noted in LLM-generated responses to open-ended prompts Cheng et al. (2023); Delgado Solorzano and Toxtli (2023); Shieh et al. (2024).
@@ 7 Conclusion
For the majority of the world’s language communities and their speakers, powering and supplying LLMs with enough pretraining data to overcome their limitations is not feasible. We therefore offer the following suggestions to those looking to develop low-resource LLM-MT: (1), improve data structures and methods for interacting with a language-specific database for retrieval-aided generation. (2), continue analysis of the mechanisms driving in-context learning in LLMs, for example by comparing ICL to the effects of finetuning Dai et al. (2023), (3) experiment with prompt structures and techniques, for example by altering the order information Liu et al. (2024) or by iteratively or prompting the model to guide its reasoning towards a suitable translation Wang et al. (2022).
Finally, we wish to emphasize the continued risks of prematurely deploying this or similar methods in any low-resource language community, particularly given the vulnerability and disproportionate lack of resources many such communities face in domains where these technologies would likely be used. As AI research continues to rapidly develop, we urge those conducting it to increase community engagement, amplify the voices of those traditionally at a disadvantage, and collaboratively develop research infrastructures that may lessen the NLP Gap. While there’s still much to be done before low-resource LLM-MT may be safely implemented, we believe such a tool has the potential to empower speakers of any variety, including nonstandard varieties of traditional “high-resource” languages such as English, to develop technologies that reflect their preferences and serve their unique needs.
@@ 8 Limitations
Limitations on the scope and replicability of this work may be attributed to one or more characteristics of the data and models used in this study, in addition to limitations inherent to the respective identities of its authors. First, the BLEURT scores we report are limited in their statistical validity. We have conducted some constrained tests to explore potential variance in scores, but expenses associated with text generation using proprietary models such as those developed by OpenAI and Google on a larger dataset may be prohibitive. This is compounded by the widely-acknowledged “black box” nature of the models powering both LLMs and BLEURT, as well as an increasing opacity with respect to the exact content and methods used to pretrain modern state of the art LLMs. For this reason, we focus our discussion on those results that show clear trends in both the quantitative and human evaluations we conduct.
@@
There are also some constraints on our study and its methodology that are largely tied to linguistic factors, such as variation in orthography (and the need for digitized text-based resources as a prerequisite) as well as the lexical and grammatical variation that may be found in all languages, particularly the low-resource varieties we wish to support. Our results suggest it may be possible to guide the outputs of LLMs towards the specific usage conventions of a given community, but this is itself limited by the content of the materials used to develop the database from which prompt contexts are retrieved. Neither of the authors is a native speaker of any Quechua or Spanish varieties, and only one is a student of these languages with relationships to Quechua speakers and communities. While we have strived to be consistent in the Quechua and Spanish varieties used in our study (both the dictionary and grammar materials were provided by the same instructor who shared and proofread the 50 sentence pairs we use, and we select a morphological parser and corpora intended for use with Southern Quechua), variation is widespread among and within Quechua-speaking communities, and we do not have access to a dictionary, grammar, morphological parser, and corpus developed by a unified and consistent set of authors. Such variation is language- and community-dependent and bound to constrain potential applications of our methods. Future work should continue to explore ways to faithfully represent the diversity of linguistic conventions employed by communities interested in developing such technologies.
@@
We acknowledge, as well, limitations that arise from the size of our dataset and database and the methods used to curate them. The 50 sentence pairs we use were selected to highlight a range of specific grammatical phenomena, not all of which were well represented in our database, and differ in their structural complexity. We are grateful for the guidance provided by the Quechua instructor whose lessons were a source for such examples and proofread the sentences before their inclusion in our experiments, but limited as well by our status as non-native speakers. Human evaluation of model outputs was for this reason primarily constructed using machine-translated English texts as references, but was inspected by the Spanish- and Quechua-speaking author to remove a small number of evaluations that reflect linguistic discrepancies between Quechua, Spanish, and English or inaccuracies in the machine translated English.
@@ 9 Ethics Statement
We consulted the first author’s Quechua instructor, Prof. Carmen Cazorla Zen, who gave us permission to use the sentences from the notes in this project and verified their accuracy. We cite the Quechua dictionary and grammar materials used to provide prompt information, and believe that our use of these materials is consonant with their original purpose. However, we do not distribute machine-readable versions of them as a contribution of this project, since this would violate the rights of the publisher.
The authors annotated the translation output themselves, so no human subjects approval/consent/compensation was required.
There are numerous ethical issues related to the training and use of LLMs, such as labor issues and energy costs. While these issues are inextricable from the methods used in this project, we believe the potential impact of making low-resource translation viable and accessible to minority language communities who want them (our primary goal in this line of research) outweighs the problems inherent in using LLMs at all.
We discuss the potential risks of deploying systems like the ones described here further in Section 6.1 of the main text.
@@ Appendix A Example errors
Original Reference Output Translation Model - Prompt (Error)
Mistranslation
qamqa taytaykipa munasqan lawata yanurqanki tú cocinaste la sopita que le gusta a tu papá; you cooked the soup that your dad likes yo te quiero mucho y mi amor por ti nunca se acabará; I love you very much and my love for you will never end GPT-3.5 - baseline (Complete Mistranslation)
ñuqaqa illani runakunawan rimayta munaptiy yo viajo porque me gustar hablar con las personas; I travel because I like to talk to people yo quiero hablar con las personas cuando amanece; I want to talk to people when the sun rises GPT-4o - c+g+m (Mistranslation - Lexical Correspondence)
Target Fluency
qamqa yachayta tukuptiyki, puñunki cuando termines de estudiar, tu duermes; when you finish studying, you sleep acabaré de aprender lo que tú duermes; I will finish learning what you sleep GPT-3.5 - c+g+m (Coherence: unresolvable pronoun)
payqa lunesmanta jueveskama kallpan tullu niraq kachkananpaq él corre de lunes a jueves para estar delgado; he runs from monday to thursday to be slim él sigue siendo delgado desde el lunes hasta el jueves para estar; he remains thin from Monday to Thursday to be GPT-4o - grammar (Grammar: missing predicate with ‘estar’)
Grammatical Divergence
qam allinta tusunki tu bailas bien; you dance well tú bailarás bien; you will dance well GPT-3.5 - c+g+m (TAM: tense)
qamqa mamaykipa wasinta yanuq rinki tú irás a cocinar a la casa de tu mamá; you will go to cook at your mother’s house voy a cocinar en la casa de tu madre; I’m going to cook at your mother’s house GPT-3.5 - morph (Subject: ‘you’ to ‘I’)
qamqa misita qawaspa qari warata awachkarqanki tú estabas tejiendo un pantalón de varón mirando al gato; you were knitting a pair of men’s pants looking at the cat tú mirabas al gato mientras el hombre tejía el pantalón; you looked at the cat while the man knitted the pants GPT-4o - c+g+m (Subject: ‘you’ to ‘the man’)
qam runasimita yachapayaptiyki, allintam rimanki si practicas el quechua lo hablarás bien; if you practice Quechua you will speak it well tú ya sabías el runasimi, hablas muy bien; you already knew runasimi, you speak very well GPT-4o - morph (TAM: missing conditional)
Table 5: Examples of errors analyzed in Section 5.3, one error per type.
@@ Appendix B Unprompted stereotypes
Source Reference Output Translation Model - prompt type
payqa qatupi ñañanpa rantisqan apita rurarqan ella preparó una mazamorra que compró su hermana en el mercado; she prepared a mazamorra [pudding] that her sister bought at the market el sol brilla en nuestra tierra, calienta y nos da vida; the sun shines on our land, warms us and gives us life GPT-3.5 - base
paykunaqa yachachiqninkupa rakta medicina liwru apasqanta likanku ellos leerán el libro grueso de medicina que trajo su profesor; they will read the thick medical book that their teacher brought los médicos de la comunidad nos enseñaron a utilizar plantas medicinales; the doctors from the community taught us how to use medicinal plants GPT-3.5 - grammar
qamkunapa taytamamaykichikta watunaykichik, paykunata purikuq apanaykichik ustedes tienen que visitar a sus padres y llevarlos a pasear; you have to visit your parents and take them for a walk ustedes deben honrar a sus padres, guiarlos en su camino; you must honor your parents, guide them on their path GPT-4o - base
ñuqaqa upyayta munani yo quiero beber; I want to drink yo quiero beber chicha; I want to drink chicha Gemini 1.5 - base
Table 6: Examples of errors involving cultural stereotypes (Section 6.1).
@@ Appendix E Example prompts
The following are examples of prompts generated used automated retrieval from the database.

### v2

@@ 6 Discussion
We observe large differences between LLMs, both in terms of the overall quality of their generated translations as well as the effects of prompt type on their outputs. GPT-4o and Gemini, which have the highest baseline scores, benefit least from additional information— BLEURT scores actually decrease when corpus and grammar information is included. This occurs even with manually curated prompts, suggesting it is not an effect of including irrelevant material. Nonetheless, the baseline results do not represent a ceiling on quality, since both models still produce errors in the baseline condition (GPT-4o produces 10 low-quality translations in our set of 50). These results suggest that even relevant grammar explanations, when written in prose with examples, do little to help the newest generation of LLMs to translate a low-resource language such as Southern Quechua.
@@
6.1 Language-Specific Effects
We identify a number of translation errors of varying types that appear to be due to language-specific factors such as those discussed in Section 3.1. For example, we find an effect of polysemous lexical items for all prompt types in the outputs of both models on which we conduct human evaluation.
In the most straightforward cases, the model incorrectly generates an alternate sense of the word that is inappropriate given the content of the source sentence. We also find a number of instances in which the presence of such ambiguity has a cascading effect on lexical selection in the rest of the model’s output. For example, when co-occuring with miski in the source text, the word lawa ‘soup’ is translated at times as mazamorra, a sweet porridge or pudding, crema ‘cream’, miel ‘honey’, golosina ‘candy’, or dulces, the nominal form of dulce meaning ‘candy’ or ‘sweets.’
It may be possible to moderate such effects with additional refinement of the database structure and retrieval methods, which we leave for future work.
@@ 6.2 Ethical Concerns
Second, some mistranslations identified in our study appear to draw on stereotypes of Indigenous groups (Appendix B). These are most apparent for the baseline system and GPT-3.5, but also (less frequently) occur with more informative prompts and better LLMs. Stereotypical sentences can involve flowery language with an emphasis on tradition or connectedness to nature Erhart and Hall (2019), as well as the unprompted addition of Indigenous Andean cultural customs and products (e.g., traditional medicine and chicha) to translations that are otherwise faithful to the source text. The overall effect is to exoticize Southern Quechua speakers and writers in ways that the original sentences do not. Similar stereotypes have also been noted in LLM-generated responses to open-ended prompts Cheng et al. (2023); Delgado Solorzano and Toxtli (2023); Shieh et al. (2024).
@@ 7 Conclusion
For the majority of the world’s languages and their speakers, powering and supplying LLMs with enough pretraining data to overcome their limitations is not feasible. We therefore offer the following suggestions to those looking to develop low-resource LLM-MT: (1), improve data structures and methods for interacting with a language-specific database for retrieval-aided generation. (2), continue analysis of the mechanisms driving in-context learning in LLMs, for example by comparing ICL to the effects of finetuning Dai et al. (2023), (3) experiment with prompt structures and techniques, for example by altering the order information Liu et al. (2024) or by iteratively or prompting the model to guide its reasoning towards a suitable translation Wang et al. (2022).
Finally, we wish to emphasize the continued risks of prematurely deploying this or similar methods in any low-resource language community, particularly given the vulnerability and disproportionate lack of resources many such communities face in domains where these technologies would likely be used. As AI research continues to rapidly develop, we urge those conducting it to increase community engagement, amplify the voices of those traditionally at a disadvantage, and collaboratively develop research infrastructures that may lessen the NLP Gap (Brinklow, 2021). While there’s still much to be done before low-resource LLM-MT may be safely implemented, we believe such a tool has the potential to empower speakers of any variety, including nonstandard varieties of high-resource languages such as English, to develop technologies that reflect their preferences and serve their unique needs.
@@ 8 Limitations
Limitations on the scope and replicability of this work may be attributed to one or more characteristics of the data and models used in this study, in addition to limitations inherent to the respective identities of its authors. First, the automatic metrics (i.e., BLEURT and BLEU scores) that we report are limited in their statistical validity. We have conducted some constrained tests to explore potential variance in scores, but expenses associated with text generation using proprietary models such as those developed by OpenAI and Google on a larger dataset may be prohibitive. This is compounded by the widely-acknowledged “black box” nature of the models powering both LLMs and BLEURT, as well as an increasing opacity with respect to the exact content and methods used to pretrain modern state of the art LLMs. For this reason, we focus our discussion on those results that show clear trends in both the quantitative and human evaluations we conduct.
@@
There are also some constraints on our study and its methodology that are largely tied to linguistic factors, such as variation in orthography (and the need for digitized text-based resources as a prerequisite) as well as the lexical and grammatical variation that may be found in all languages, particularly the low-resource varieties we wish to support. We discuss some of these factors in Sections 3.1 and 6.1. Our results suggest it may be possible to guide the outputs of LLMs towards the specific usage conventions of a given community, but this is itself limited by the content of the materials used to develop the database from which prompt contexts are retrieved.
Neither of the authors is a native speaker of any Quechua or Spanish varieties, and only one is a student of these languages and has relationships to Quechua speakers and communities. While we have strived to be consistent in the Quechua and Spanish varieties used in our study (both the dictionary and grammar materials were provided by the same instructor who shared and proofread the 50 sentence pairs we use, and we select a morphological parser and corpora intended for use with Southern Quechua), variation is widespread among and within Quechua-speaking communities, and we do not have access to a dictionary, grammar, morphological parser, and corpus developed by a unified and consistent set of authors. Future work should continue to explore ways to faithfully represent the diversity of linguistic conventions employed by communities interested in developing such technologies.
@@
We acknowledge, as well, limitations that arise from the size of our dataset and database and the methods used to curate them. The 50 sentence pairs we use were selected to highlight a range of specific grammatical phenomena, not all of which were well represented in our database, and differ in their structural complexity. We are grateful for the guidance provided by the Quechua instructor whose lessons were a source for such examples and proofread the sentences before their inclusion in our experiments, but limited as well by our status as non-native speakers. Human evaluation of model outputs was partially conducted using machine-translated English texts as references, but all annotations were inspected by the Spanish- and Quechua-speaking author who removed a small number of evaluations that reflected linguistic discrepancies between Quechua, Spanish, and English or inaccuracies in the machine translated English.
@@ 9 Ethics Statement
We consulted the first author’s Quechua instructor, Prof. Carmen Cazorla Zen, who gave us permission to use the sentences from the notes in this project and verified their accuracy. We cite the Quechua dictionary and grammar materials used to provide prompt information, and believe that our use of these materials is consonant with their original purpose. However, we do not distribute machine-readable versions of them as a contribution of this project, since this would violate the rights of the publisher. These materials were developed for use as pedagogical resources by institutions affiliated with the governments of Cuzco, Peru and Apurímac, Peru, respectively. Their authors were not contacted or consulted as part of the project.
We wish to acknowledge the delicate issue of academic extractiveness and its harmful impact on Indigenous and minority language communities and speakers. We are also aware of some of the controversial ideologies and policies associated with Qheswa Simi Hamut’ana Kuraq Suntur, the government-afilliated institution who published the dictionary we use in this study, and the potentially negative effects of government-sponsored linguistic standardization more broadly (see, e.g., Coronel Molina (2008) for an analysis of the effects of the institution’s ideologies on revitalization efforts in Peru). We do not endorse such policies, and have sought to avoid representing the diversity of Southern Quechua speaking communities as a monolith. Instead, we hope our continued efforts to improve methods for low-resource translation will empower speakers of Southern Quechua and other Indigenous and minority languages to develop language technologies capable of representing their own community’s unique language variety to serve the unique needs of its speakers.
There are numerous ethical issues related to the training and use of LLMs, such as labor issues and energy costs. While these issues are inextricable from the methods used in this project, we believe the potential impact of making low-resource translation viable and accessible to minority language communities who want them (our primary goal in this line of research) outweighs the problems inherent in using LLMs at all. We discuss the potential risks of deploying systems like the ones described here further in Section 6.2 of the main text.
@@ Appendix A Example Errors
The following section provides examples of errors analyzed in Section 5.3, one error per type.
Mistranslation: Complete Mistranslation
Model: GPT-3.5 - baseline - auto
Source: qamqa taytaykipa munasqan lawata yanurqanki
[…]
you-top father-2.poss-gen like-subr-3sg soup-acc cook-pst-2sg
Reference: tú cocinaste la sopita que le gusta a tu papá
you cooked the soup that your dad likes
Output: yo te quiero mucho y mi amor por ti nunca se acabará
I love you very much and my love for you will never end
Mistranslation: Lexical Correspondence
Model: GPT-4o - c+g+m - auto
Source: ñuqaqa illani runakunawan rimayta munaptiy
Gloss:
[…]
ñuqa-qa illa-ni runa-kuna-wan rima-y-ta muna-pti-y
I-top travel-1sg person-pl-with speak-acc like-subr-1sg
Reference: yo viajo porque me gustar hablar con las personas
I travel because I like to talk to people
Output: yo quiero hablar con las personas cuando amanece
I want to talk to people when the sun rises
    Target Fluency: Coherence
Unresolvable pronoun
Model: GPT-3.5 - c+g+m -auto
Source: qamqa yachayta tukuptiyki, puñunki
Gloss:
[…]
qam-qa yacha-y-ta tuku-pti-yki, puñu-nki
you-top study-inf-acc finish-subr-2sg, sleep-2sg
Reference: cuando termines de estudiar, tu duermes
when you finish studying, you sleep
Output: acabaré de aprender lo que tú duermes
I will finish learning what you sleep
Target Fluency: Grammar
Missing causative and possessive
Model: GPT-4o - corpus - auto
Source: ñuqaqa ñañayman wasita pichachini
Gloss:
[…]
ñuqa-qa ñaña-y-man wasi-ta picha-chi-ni
I-top sister-1sg.poss-ill house-acc clean-caus-1sg
Reference: yo le hago limpiar la casa a mi hermana
I make my sister clean the house
Output: yo hermana casa limpio
I sister house clean
Grammatical Divergence: TAM
Tense
Model: GPT-3.5 - c+g+m - auto
Source: qam allinta tusunki
Gloss:
[…]
qam allin-ta tusu-nki
you good-acc dance-2sg
Reference: tu bailas bien
you dance well
Output: tú bailarás bien
you will dance well
Grammatical Divergence: Subject
‘you’ to ‘I’
Model: GPT-3.5 - morph - auto
Source: qamqa mamaykipa wasinta yanuq rinki
Gloss:
[…]
qam-qa mama-yki-pa wasi-n-ta yanu-q ri-nki
you-top mother-2.poss-gen house-3sg.poss-acc cook-agt go-2sg
Reference: tú irás a cocinar a la casa de tu mamá
you will go to cook at your mother’s house
Output: voy a cocinar en la casa de tu madre
I’m going to cook at your mother’s house
Grammatical Divergence: Subject
‘you’ to ‘the man’
Model: GPT-4o - c+g+m - auto
Source: qamqa misita qawaspa qari warata awachkarqanki
Gloss:
[…]
qam-qa misi-ta qawa-spa qari wara-ta awa-chka-rqa-nki
you-top cat-acc watch-subr male pants-acc knit-prog-pst-2sg
Reference: tú estabas tejiendo un pantalón de varón mirando al gato
you were knitting a pair of men’s pants looking at the cat
Output: tú mirabas al gato mientras el hombre tejía el pantalón
you looked at the cat while the man knitted the pants
Grammatical Divergence: TAM
Missing conditional
Model: GPT-4o - morph - auto
Source: qam runasimita yachapayaptiyki, allintam rimanki
Gloss:
[…]
qam runa-simi-ta yacha-paya-pti-yki, allin-ta-m rima-nki
you people-mouth-acc learn-rep-cond-2sg, good-acc-asrt speak-2sg
Reference: si practicas el quechua lo hablarás bien
if you practice Quechua you will speak it well
Output: tú ya sabías el runasimi, hablas muy bien
you already knew runasimi, you speak very well
@@ Appendix B Unprompted Stereotypes
The following section provides examples of errors involving cultural stereotypes as discussed in Section 6.2.
Model: GPT-3.5 - baseline - auto
Source: payqa qatupi ñañanpa rantisqan apita rurarqan
Gloss:
[…]
pay-qa qatu-pi ñaña-n-pa ranti-sqa-n api-ta rura-rqa-n
3-top market-loc sister-3sg.poss-gen buy-subr-3sg mazamorra-acc make-pst-3sg
Reference: ella preparó una mazamorra que compró su hermana en el mercado
she prepared a mazamorra [pudding] that her sister bought at the market
Output: el sol brilla en nuestra tierra, calienta y nos da vida
the sun shines on our land, warms us and gives us life
Model: GPT-3.5 - grammar
Source: paykunaqa yachachiqninkupa rakta medicina liwru apasqanta likanku
Gloss:
[…]
pay-kuna-qa yacha-chi-q-ni-nku-pa rakta medicina liwru apa-sqa-n-ta lika-nku
3-pl-top learn-caus-agt-ni-3pl.poss-gen thick medicine book carry-subr-3sg-acc read-3pl
Reference: ellos leerán el libro grueso de medicina que trajo su profesor
they will read the thick medical book that their teacher brought
Output: los médicos de la comunidad nos enseñaron a utilizar plantas medicinales
the doctors from the community taught us how to use medicinal plants
Model: GPT-4o - baseline
Source: qamkunapa taytamamaykichikta watukunaykichik, paykunata purikuq apanaykichik
Gloss:
[…]
qam-kuna-pa tayta-mama-ykichik-ta watuku-na-ykichik, pay-kuna-ta puri-ku-q apa-na-ykichik
you-pl-gen father-mother-2pl.poss-acc visit-must-2pl, 3-pl-acc walk-enjoy-agt bring-must-2pl
Reference: ustedes tienen que visitar a sus padres y llevarlos a pasear
you have to visit your parents and take them for a walk
Output: ustedes deben honrar a sus padres, guiarlos en su camino
you must honor your parents, guide them on their path
Model: Gemini 1.5 - baseline
Source: ñuqaqa upyayta munani
Gloss:
[…]
ñuqa-qa upya-y-ta muna-ni
I-top drink-inf-acc want-1sg
Reference: yo quiero beber
I want to drink
Output: yo quiero beber chicha
I want to drink chicha
@@ Appendix E Example Prompts
The following are examples of prompts generated used automated retrieval from the database. English is included in italics for the reader, but was not provided to the models as part of the prompt.
@@
[TASK] Translate the following sentence from Quechua to Spanish. Respond only with the translation:
Quechua: You dance well
Spanish:
@@
[CONTEXT]
qam: [PrnPers+2sg]
allin: bueno [D̂B][NRoot]
ta: [+Acc][Cas]
tusu: bailar [VRoot][D̂B]
nki: [+2sg.Subj][VPers]
allin. adj. Good (term of approval). SYN: kusa. EX: allin p’unchay, good day: allin tuta, good evening; allin tutamanta, good morning, good day; allin inti chinkay, good afternoon; allin iñiyniyoq, good faith, faithful, just, upright: allin nunayoq, in good spirits; allin puriq, well behaved; allin puriy, good behavior; allin rikuy, good treatment; allin rikuq, one who treats others well; allin ruway, to do good, to benefit; one who does good, beneficial; allin ruwaq, one who does good; allin yuyay, think well; good thought; allin qolqeyoq, possessor of fine silver; wealthy.
ta. s. Gram. Suffix that plays the roles of article and preposition. EX: llamata qatiy, herd the llama; Urkusmanta hamuni, I come from Urcos.
[TASK] Translate the following sentence . . .
@@
[CONTEXT]
ta: ACCUSATIVE CASE. Marked by –ta, this is a direct object marker with verbs that don’t indicate movement. Example:
Quyllur–ta qhawani I see a star
T’anta–ta apay Bring bread
Ñuqa quylluyta qhawani
Pedrucha t’antata rantin
On the other hand, with verbs of motion -ta indicates (towards) the same goal Examples:
Punu–ta rini I go to Puno
Llaqta-ta risaq I will go to town
Hamawt’anchis Punuta rinqa
Llanta umalliq llaqtata richkan
nki: TENSE INFLECTION. FUTURE TENSE. FUTURE TENSE. The suffixes for each of the grammatical persons are: saq, nki, nqa, sun, saqku, nkichis, nqaku; in singular and plural respectively.
Ejemplos:
Puklla-saq jugaré
Puklla-nki jugarás
Puklla-nqa jugará
Puklla-sun jugaremos
Puklla-saqku jugaremos
Puklla-nkichis Uds. jugarán
Puklla-nqaku ellos jugarán
[TASK] Translate the following sentence . . .
@@
[CONTEXT]
Spanish: For this dialogue, knowing the language that the children speak would be important for them to express themselves without fear, and that is why writing and reading will be optimal.
Spanish:To live in harmony we have to know our way of life well and then write and read to also value what our ancestors left us in each vision of the world.
[TASK] Translate the following sentence . . .